A production process consists of a three-step operation. The scrap rate is 13 percent for the first step and 8 percent for the other two steps.

a.If the desired daily output is 488 units, how many units must be started to allow for loss due to scrap? (Do not round intermediate calculations. Round up your final answer to the next whole number.)

Number of units            

b.If the scrap rate for each step could be cut in half at every operation, how many units would this save in terms of the scrap allowance? (Do not round intermediate calculations. Round up your final answer to the next whole number.)

Number of units            

c.If the scrap represents a cost of $10 per unit, how much is it costing the company per day for the original scrap rate (i.e. the Part a scrap rate)? (Round your final answer to the nearest whole number. Omit the "$" sign in your response.)

Cost           $  
Answer #1

(a)

Step Input Output
1 662.712 576.560
2 576.560 530.435
3 530.435 488.000

So, no. of units with which to be started = 663

(b)

Step Input Output
1 566.325 529.514
2 529.514 508.333
3 508.333 488.000

Savings in terms of scrap units = 662.712 - 566.325 = 97 units (or, 96, depends on rounding off)

(c)

Scrap cost per day = (663 - 488) * $10 = $1,750